A projectile is fired from the ground with an upward speed of 245 m/s. After how many seconds will it return to the ground? Use the formula h=vt-4.9t^2
Answer by Alan3354(69443)

A projectile is fired from the ground with an upward speed of 245 m/s. After how many seconds will it return to the ground? Use the formula h=vt-4.9t^2
---------------------
vt - 4.9t^2 = 0
245t - 4.9t^2 = 0
t*(245 - 4.9t) = 0
t = 0 (when it's fired)
t = 50 seconds (when it returns)